OpenStudy (anonymous):

Solve for y: 2/3y + 4 = 6

OpenStudy (anonymous):

y = 3

OpenStudy (anonymous):

2/3y = 2 y = 1/3 y = 1/3

OpenStudy (anonymous):

2 --y = 2 3

OpenStudy (anonymous):

Medals to everyone!

OpenStudy (anonymous):

y=3